# Neural Conditional Transport Maps

### v1

@@ 3.5 Training Procedure
4 Results
In this section, we present a comprehensive empirical evaluation of our framework. We begin by describing our datasets, which include real scientific data from climate-economy and Integrated Assessment Models. Then, we will present the results of our ablation studies, first examining the impact of simple but effective improvements with respect to the unconditional formulation in Korotin et al. (2023), then systematically evaluating different design choices on the conditional setting, including the use of pretraining and the type of conditioning. Finally, we will show results on the benchmarking tasks and discuss computational aspects of them. Benchmarking data will be released upon publication, more results are provided in the supplementary materials.
4.1 Applications
Climate Economic Impact Distribution Transport
We examine the economic impacts of climate change using the empirical damage function model of Burke et al. (2015). This application requires building an emulator for complex multivariate distributions conditioned on categorical (scenario) and continuous (time) variables.
Problem Formulation. Let $\mathcal{X}\subset\mathbb{R}^{n}$ represent the space of GDP per capita with climate damages across $n=20$ countries. For each country $i$, the impact is quantified through 1000 bootstrap replicates, resulting in a high-dimensional empirical distribution. We define our conditioning space $\mathcal{C}=\mathcal{C}_{\text{ssp}}\times\mathcal{C}_{\text{year}}$, where $\mathcal{C}_{\text{ssp}}=\{0,1,2,3\}$ represents four SSP scenarios (SSP1-1.9, SSP2-4.5, SSP3-7.0, SSP5-8.5) and $\mathcal{C}_{\text{year}}=[2030,2100]$ represents projection years. Our goal is to learn a conditional transport map $T_{\theta}:\mathbb{R}^{n}\times\mathcal{Z}\times\mathcal{C}\rightarrow\mathcal{X}$ that efficiently transforms samples from a reference distribution to match target distributions under different climate scenarios and future years. Thus, we enable efficient sampling of climate impact distributions while preserving their statistical properties, which is needed for appropriate uncertainty quantification.
Global Sensitivity Analysis for Integrated Assessment Models
Our second application focuses on global sensitivity analysis (GSA) for the RICE50+ IAM (Gazzotti, 2022), using the OT-based sensitivity indices presented in Borgonovo et al. (2024).
Problem Formulation. Let $f:\mathcal{C}\subset\mathbb{R}^{3}\rightarrow\mathcal{Y}\subset\mathbb{R}^{58}$ be the RICE50+ model mapping input parameters $\mathbf{C}\in\mathcal{C}$ to the output $\mathbf{Y}\in\mathcal{Y}$. RICE50+  (Gazzotti, 2022) is an IAM with high regional heterogeneity used to assess climate policy benefits and costs. We estimate the sensitivity of the CO2 emissions for a single region (the output $\mathbf{Y}$) to three inputs $\mathbf{C}$, related to the emissions abatement costs (klogistic), the aversion to inter-country inequality (gamma), and the climate impacts (kw_2), leveraging a subset of the data available in Chiani et al. (2025). Following Borgonovo et al. (2024), the OT-based sensitivity index for variable $C_{i}$ is defined as: $\iota^{K}(\mathbf{Y},C_{i})=\frac{\mathbb{E}_{C_{i}}[K(\mu_{\mathbf{Y}},\mu_{\mathbf{Y}|C_{i}})]}{\mathbb{E}[k(\mathbf{Y},\mathbf{Y}^{\prime})]}$, where $\mu_{\mathbf{Y}}$ is the unconditional output distribution, $\mu_{\mathbf{Y}|C_{i}}$ is the conditional output distribution when variable $C_{i}$ is fixed, and $K$ is the OT cost defined in Eq. (1). Computing these indices traditionally requires solving multiple OT problems for each conditioning variable and partition, which becomes computationally intractable for high-dimensional models or large datasets. Instead, we partition each input space into $M=25$ bins and define our conditioning space $\mathcal{\tilde{C}}=\{0,1,2\}\times[0,1]$, where the first component represents the variable index and the second represents the normalised partition. Our approach learns a single conditional transport map $T_{\theta}:\mathcal{Y}\times\mathcal{Z}\times\mathcal{\tilde{C}}\rightarrow\mathcal{Y}$ parameterised by neural networks.
@@ 4.1 Applications
Ours Pretrain Embed. Conditioning Type Continuous Encoding
@@
False Shared Concat SE FiLM FAN Attn. CLN AdaIN Scalar Fourier
@@
Time (s) $\downarrow$ 466 451 457 451 554 854 487 921 712 658 464 486
Parameters (k) $\downarrow$ 1158 1158 1156 1324 2764 2874 2861 2908 2791 2790 1150 1152
@@
Wass. $\downarrow$ 0.170 0.214 0.265 0.267 0.716 0.974 0.458 0.721 0.691 0.744 0.255 0.172
Wass. $\times$ time $\downarrow$ 79.22 96.51 121.1 120.4 396.7 831.8 223.0 664.0 491.9 489.6 118.3 83.59
@@
$\rho$uparrow 0.942 0.905 0.812 0.305 0.377 0.456 0.388 0.441 0.612 0.682 0.894 0.931
$\rho/\text{time}(\times 100)\uparrow$ 0.202 0.200 0.177 0.068 0.068 0.053 0.079 0.048 0.086 0.104 0.193 0.192
Table 1: Ablation study of our conditional transport framework. We report computational cost and accuracy on our datasets. Our final model (leftmost column) uses a hypernetwork with positional encoding and pretraining, without shared embedding. Each other column group represents variations from this configuration. Best results are in green bold, second best are underlined, worst are red.
4.2 Ablation Studies
Unconditional Setting. We first evaluate architectural improvements to the base NOT framework Korotin et al. (2023). We build upon their ReLU-based architectures, incrementally adding orthogonal initialisation, residual connections, and layer normalisation, while preserving identical training configuration and comparable parameter counts across models. In Figure 3, we present results across three datasets. The first two rows show 2D probability distributions, using simple MLPs as our baseline architecture. We measure the KL divergence and Wasserstein loss between targets and transported distributions (from a 2D Uniform prior), using $2^{15}$ samples. The third row showcases a toy convolutional model performing Image-to-Image translation on MNIST (digit 2 to digit 3), evaluated with the MMD distance, Wasserstein loss, FID score, and classification accuracy. In both MLP and convolutional settings, our enhancements yield notable quantitative and qualitative improvements, as shown on the right panel. These architectural enhancements are the foundation of our conditional transport experiments, improving convergence without increasing computational cost.
Conditional Transports. We present our ablation study for our conditional framework in Table 1. We compare our best model configuration, which uses hypernetwork conditioning, pretraining, positional encoding, and separate conditioning embeddings for $T$ and $f$, with variations of that model. For each, we report training time and number of parameters, as well as absolute and cost-adjusted accuracy on two datasets: Climate damages and IAMs. We use the same training setup across datasets and model configurations. Training times are reported for the climate damages dataset. For IAM, we report the Pearson’s correlation $\rho$ between the costs obtained by our model, and those computed with simplex. All experiments are run on a RTX 4070 GPU and an i7-13700H CPU. Using CodeCarbon Lacoste et al. (2019), we measure an electricity consumption of 0.009323 kWh.
Interesting patterns emerge from these results. First, we observe that the conditioning type plays a significant role in training time and accuracy. Our lightweight hypernetwork consistently outperforms simpler alternatives like feature modulation or adaptive normalization. Importantly, we achieve this without a major increase in training cost or trainable parameters. Other complex approaches like cross-attention are suboptimal in accuracy and cost. Concatenation provides the lowest training time, but achieves the worst accuracy on the IAM setting. The encoding performed on the continuous variables also significantly impacts both datasets, with Positional Encoding (our solution) and Fourier features providing the best accuracy, albeit the latter shows less efficiency. This is true for both datasets, where the continuous variable has distinct meanings, suggesting that adequate processing of the raw conditioning plays an important role in expressivity. We also tested whether to share the conditioning embeddings (with hypernetworks) for $T$ and $f$, as proposed in CondOT Bunne et al. (2022). Our results show that separating these embeddings leads to accuracy gains, suggesting that $T$ and $f$ use the condition in distinct ways. Finally, we show that our pretraining algorithm enables higher overall accuracy with no significant added training cost. These results show that adequate, data-driven initialization can improve training dynamics and overall results without efficiency losses.
4.3 Results
Climate Economic Impact Distributions: Figure 4 presents the performance of our model on the climate damages dataset, showing its capability to generate realistic distributions across different climate scenarios. For each SSP scenario, we compare ground truth GDP per capita with climate damages distributions with samples from our conditional transport model over the 2030-2100 time horizon. The results demonstrate our model’s effectiveness at capturing both central tendencies and uncertainty (shaded regions, $90\%$ confidence intervals), specific to each scenario. Our approach learns the distinct patterns across different SSPs: SSP1 shows relatively stable relative impacts with wide uncertainty, while SSP2 exhibits a moderate decline after 2060. The SSP3 scenario shows the most pronounced downward trend, which our model captures accurately despite the smaller uncertainty.
@@ 4.3 Results
Global Sensitivity Analysis: Figure 5 presents a comparison between our neural transport method and the traditional simplex-based approach for computing OT-based sensitivity indices across three input variables in the RICE50+ model. The first three panels show the transport costs across different partition values (0 to 1) for each variable, with simplex results in blue and our neural method in red. Our neural transport approach closely tracks the simplex cost patterns across all three variables, effectively capturing the complex sensitivity structure of the underlying model. For the klogistic variable (first panel), both methods identify similar regions of high sensitivity, with our neural approach maintaining strong correlation with simplex results. The gamma variable (second) shows more complex sensitivity patterns that our method accurately reproduces, including the pronounced peak near partition value 0.25. For kw_2 (third), both methods detect lower overall sensitivity with consistent patterns across partition values.
The rightmost panel summarizes the average transport costs for each variable. Our neural method preserves the relative importance ordering among variables while maintaining comparable absolute cost values (note that the magnitude of simplex costs may not be comparable with ours). This is critical for GSA applications where accurate ranking of variable importance drives decision-making. Notably, our neural approach achieves this accuracy with higher scalability—requiring a single trained model rather than solving hundreds of individual OT problems. These results demonstrate that our framework enables efficient global sensitivity analysis for complex models, opening new possibilities for comprehensive uncertainty quantification in high-dimensional modeling domains.
@@ 5 Conclusions
We have presented a neural framework for learning conditional OT maps between probability distributions that extends the NOT framework to handle both categorical and continuous conditioning variables. Our hypernetwork-based architecture generates transport layer parameters dynamically, creating adaptive mappings that significantly outperform simpler conditioning approaches across diverse benchmarks. Experiments on synthetic and real-world datasets demonstrate superior performance compared to existing methods, with ablation studies confirming the value of each architectural component. We have also shown how our approach can effectively be applied to global sensitivity analysis, offering high computational efficiency while maintaining theoretical guarantees.
Limitations and Future Work. Our approach has several limitations: we primarily evaluated residual feedforward networks, not testing recurrent architectures, we have not explored conditioning on complex modalities like CLIP Radford et al. (2021) embeddings or pixel-wise semantic labels, and our hypernetwork approach incurs higher computational costs than simpler alternatives. Besides, our method has been primarily evaluated on climate-related data. Future work could explore multi-modal conditionings, more efficient architectures, applications to image generative models, dynamical systems and causal inference, or connections to diffusion models and normalising flows.
@@
Broader Impacts. Our work could positively impact several fields by enabling more efficient uncertainty quantification and global sensitivity analysis for complex models in climate or economics, potentially leading to better-informed policy decisions. The approach also benefits controlled generative modelling applications and makes advanced optimal transport techniques more accessible to researchers with limited computational resources. However, like most ML techniques, our method could be misused if applied to sensitive domains without appropriate oversight, such as creating more realistic synthetic media. Mitigation strategies include implementing proper access controls, combining sensitivity analysis with domain expertise, and providing educational resources to help practitioners correctly interpret these methods. We believe the benefits outweigh potential risks when appropriate safeguards are implemented.
Acknowledgements
We thank Marta Mastropietro for providing us with the climate damages data. This project received funding from the European Union European Research Council (ERC) Grant project No 101044703 (EUNICE).

### v2

@@ 5.1.1 Architectural Foundation
In both MLP and convolutional settings, our enhancements yield notable quantitative and qualitative improvements, as shown in the visualization panels. These architectural enhancements form the foundation of our conditional transport experiments, demonstrating that improved convergence can be achieved without increasing computational cost.
5.1.2 Conditional Framework Design
Having established our architectural foundation, we turn to the key design decisions specific to conditional transport learning. Table 1 presents our systematic evaluation of conditioning mechanisms, encoding strategies, and training procedures.
Table 1: Ablation study of our conditional transport framework. We report computational cost and accuracy on our datasets. Our final model (leftmost column) uses a hypernetwork with positional encoding and pretraining, without shared embedding. Each other column group represents variations from this configuration. We present best results for our method in green bold, second best are underlined, worst are red. Note that for the MNIST experiment, there is no continuous input variable, therefore we don’t ablate their encodings. CondOT default uses the original CondOT architecture and optimization recipe with tokenized factorized conditioning over categorical and discretized continuous variables. CondOT adapted keeps the CondOT training recipe unchanged but replaces this discretized conditioning with mixed conditioning combining one-hot categorical inputs and normalized continuous scalars. CondOT is evaluated only on the climate and GSA/RICE tasks; MNIST entries are marked as not applicable.
@@ 5.1.2 Conditional Framework Design
CondOT(2022) Ours Pretrain Embed. Conditioning Type Continuous Encoding
@@
Default Adapted False Shared Concat SE FiLM FAN Attn. CLN AdaIN Scalar Fourier
@@
Time (s) $\downarrow$ 586 721 466 451 457 451 554 854 487 921 712 658 464 486
Parameters (k) $\downarrow$ 1216 1316 1158 1158 1156 1324 2764 2874 2861 2908 2791 2790 1150 1152
@@
Wass. $\downarrow$ 0.322 0.255 0.170 0.214 0.265 0.267 0.716 0.974 0.458 0.721 0.691 0.744 0.255 0.172
Wass. $\times$ time $\downarrow$ 188.7 183.8 79.22 96.51 121.1 120.4 396.7 831.8 223.0 664.0 491.9 489.6 118.3 83.59
@@
$\rho$ $\uparrow$ 0.025 0.389 0.942 0.905 0.812 0.305 0.377 0.456 0.388 0.441 0.612 0.682 0.894 0.931
$\rho/\text{time}(\times 100)\uparrow$ 0.004 0.054 0.202 0.200 0.177 0.068 0.068 0.053 0.079 0.048 0.086 0.104 0.193 0.192
MNIST
Time (s) $\downarrow$ N.A. 314 306 314 277 293 319 314 326 315 311 N.A.
Parameters (m) $\downarrow$ N.A. 2.893 2.893 2.803 1.815 1.855 1.964 1.832 1.866 1.832 1.965 N.A.
Accuracy $\uparrow$ N.A. 1.000 1.000 1.000 0.992 0.600 1.000 1.000 0.800 1.000 1.000 N.A.
FID $\downarrow$ N.A. 2.509 2.656 3.167 3.565 11.18 2.994 2.732 5.792 2.855 2.912 N.A.
FID $\times$ time $\downarrow$ N.A. 787.8 803.5 994.4 987.5 3245 955.1 857.8 1888 899.3 905.6 N.A.
We compare our final configuration—using hypernetwork conditioning, pretraining, positional encoding, and separate conditioning embeddings for transport and critic networks—with systematic variations. Training times and parameter counts provide computational cost analysis, while application-specific metrics evaluate transport quality. For the climate damages dataset, we report Wasserstein distance and efficiency (Wasserstein × time). For IAM data, we report Pearson correlation $\rho$ between neural and simplex-computed costs and efficiency-adjusted correlation. All experiments use identical training protocols on RTX 4070 GPU with i7-13700H CPU, measuring 0.009323 kWh electricity consumption via CodeCarbon for the climate damages application Lacoste et al. (2019).
Several important patterns emerge from these results. The conditioning mechanism choice significantly impacts both training efficiency and accuracy. Our lightweight hypernetwork consistently outperforms alternatives like feature modulation (FiLM) or adaptive normalization approaches (AdaIN, CLN) without major increases in training cost or parameters. Complex approaches like cross-attention prove suboptimal in both accuracy and computational cost. While concatenation offers lowest training time, it achieves worst accuracy on the IAM dataset, confirming our hypothesis that simple conditioning fails when conditions require fundamentally different transport behaviors. Interestingly, adaptive normalization mechanisms like FAN, AdaIn or CLN are good alternatives to the hypernetwork in image generation tasks, which is perhaps unsurprising given their success in different computer vision problems. This also suggests that these alternatives would scale consistently better than cross-attention mechanisms, achieving both relatively expressive transformations while having a reduced training cost.
The continuous variable encoding strategy also proves crucial across both datasets. Positional encoding (our approach) and Fourier features achieve the best accuracy, though Fourier features are less computationally efficient. This holds across both applications despite different continuous variable semantics, suggesting that appropriate processing of raw conditioning information plays a vital role in expressivity.
Our analysis of embedding sharing strategies shows that separating conditioning embeddings for transport and critic networks leads to accuracy gains. This suggests that transport and critic functions utilize conditioning information in fundamentally different ways. This has important implications for conditional transport architecture design. Finally, our pretraining procedure enables higher accuracy with minimal additional training cost, demonstrating that appropriate data-driven initialization improves training dynamics and final results without efficiency penalties.
Comparisons with previous work When comparing our framework against the CondOT baselines, the results in 1 show improvements in both expressivity and computational efficiency. On the climate damages dataset, our method outperforms both CondOT Default and CondOT Adapted. In the GSA application, CondOT Default fails to capture the underlying sensitivity structures ($\rho=0.025$), and while the Adapted version improves this ($\rho=0.389$), both fall short of our hypernetwork approach ($\rho=0.942$). These results show that both adequate conditioning mechanisms designs and removing the requirements of partial convexity from neural OT architectures allow for more expressive models without increasing overall computational cost.
5.2 Application Results
Building on the insights from our ablation studies, we now demonstrate the practical effectiveness of our approach across the scientific and computational applications described in Section 4.
5.2.1 Climate Economic Impact Distributions
Figure 4 presents our model’s performance on the climate damages dataset, demonstrating its capability to generate realistic distributions across different climate scenarios. For each SSP scenario, we compare ground truth GDP per capita with climate damages distributions against samples from our conditional transport model over the 2030-2100 time horizon.
@@ 5.2.1 Climate Economic Impact Distributions
The results demonstrate our model’s effectiveness at capturing both central tendencies and uncertainty (shaded regions represent 90% confidence intervals) specific to each scenario. Our approach successfully learns the distinct temporal evolution patterns across different SSPs: SSP1 shows relatively stable relative impacts with wide uncertainty reflecting the optimistic but uncertain nature of this scenario, while SSP2 exhibits a moderate decline after 2060 consistent with intermediate climate projections. The SSP3 scenario shows the most pronounced downward trend—representing severe climate impacts—which our model captures accurately despite the smaller uncertainty bounds characteristic of this more deterministic high-impact scenario.
5.2.2 Global Sensitivity Analysis
Figure 5 presents our validation against traditional simplex-based approaches for computing OT-based sensitivity indices across three input variables in the RICE50+ model. This comparison directly tests our framework’s ability to replace computationally intensive traditional methods with efficient neural approximations.
@@ 5.2.2 Global Sensitivity Analysis
The first three panels show transport costs across different partition values (0 to 1) for each variable, with simplex results in blue and our neural method in red. Our neural transport approach closely tracks the simplex cost patterns across all three variables, effectively capturing the complex sensitivity structure of the underlying model. For the klogistic variable (emissions abatement costs), both methods identify similar regions of high sensitivity, with our neural approach maintaining strong correlation ($\rho=0.942$ from Table 1) with simplex results. The gamma variable (inequality aversion) shows more complex sensitivity patterns that our method accurately reproduces, including the pronounced peak near partition value 0.25. For kw_2 (climate impacts), both methods detect lower overall sensitivity with consistent patterns across partition values.
The rightmost panel summarizes average transport costs for each variable, revealing that our neural method preserves the relative importance ordering among variables while maintaining comparable absolute cost magnitudes. This preservation of variable ranking is critical for GSA applications where accurate identification of the most influential parameters drives decision-making in policy and scientific contexts.
Crucially, our neural approach achieves this accuracy with dramatically improved scalability—requiring a single trained model rather than solving hundreds of individual OT problems (75 separate optimizations for this three-variable, 25-partition case). This computational advantage becomes increasingly important as model complexity and conditioning dimensionality grow, enabling sensitivity analysis for problems previously considered computationally intractable.
5.2.3 Image Generation
Figure 6 presents qualitative results of our ablation study, on MNIST image generation task, using our Conditional Optimal Transport framework. As shown, our baseline configuration (top row) achieves clean digits, significantly outperforming other conditioning methods. Removing the pre-training step or sharing a the conditioning between $T$ and $f$ yield small but noticeable artifacts in the generated digits. Note that MNIST generation is a relatively trivial task for image generative models, and these networks can be trained in approximately 5 minutes on a standard laptop. How these results would scale up to large-scale image generation is outside the scope of this work, as Neural Optimal Transport is not a state-of-the-art formulation for high-resolution controllable image generation, where diffusion models or large-scale generative adversarial networks are currently among the best-performing alternatives.
Figure 6: Qualitative ablation study on MNIST digit generation, comparing our method with a hypernetwork (top row) with ablated components, including the lack of pretraining (second row), the used of shared conditioning for $T$ and $f$ (third), and the different conditioning types we tested.
5.3 Computational Analysis and Discussion
The experimental results demonstrate that our conditional neural transport framework successfully addresses the key limitations identified in existing approaches while providing practical advantages for real-world applications.
Our ablation studies reveal that the hypernetwork conditioning mechanism provides the best balance of expressivity and computational efficiency. Unlike feature modulation approaches that can only adjust existing computations, hypernetworks generate fundamentally different transformation parameters for each condition—essential when optimal transport maps must adapt qualitatively across conditioning values. The computational overhead of our lightweight hypernetwork implementation proves minimal compared to the expressivity gains achieved.
The importance of appropriate continuous variable encoding, demonstrated consistently across both climate and GSA applications, highlights an often-overlooked aspect of conditional transport learning. Our sinusoidal positional encoding effectively captures multi-scale variation in continuous conditioning variables, enabling smooth interpolation between conditioning values while maintaining sufficient expressiveness for complex conditional relationships.
The success of our pretraining strategy confirms the importance of initialization in conditional transport learning. By starting transport networks near identity mappings and stabilizing critic functions through multi-objective pretraining, we avoid the optimization pathologies that plague conditional transport training, particularly when handling diverse conditioning values that may require very different optimal mappings.
For practical applications, our approach enables new possibilities in scientific computing and uncertainty quantification. The climate modeling results demonstrate effective emulation of complex multivariate distributions under hybrid conditioning. At the same time, the GSA application shows how computational bottlenecks in sensitivity analysis can be overcome through efficient neural approximation. The preserved statistical properties and correlation with traditional methods indicate that our neural approach can be a reliable replacement for classical techniques in appropriate contexts.
Future work should explore extensions to higher-dimensional conditioning spaces and investigate theoretical guarantees for conditional transport learning. The framework’s modular design makes it well-suited for adaptation to emerging applications in generative modeling, domain adaptation, and scientific computing, where conditional optimal transport principles provide natural problem formulations.
@@ 6 Conclusions
We presented a neural framework for conditional optimal transport that addresses fundamental limitations in existing approaches. Our hypernetwork-based conditioning mechanism addresses three critical bottlenecks: the computational explosion of solving hundreds of individual OT problems in applications like global sensitivity analysis, the architectural constraints of ICNN-based methods that limit expressivity, and the inability of simple conditioning strategies to expressively model transport behaviors across conditions. Our hypernetwork architecture generates transport layer parameters dynamically based on conditioning inputs. This provides the expressiveness of separate networks per condition while maintaining computational efficiency. Our design enables different mappings for different conditions, essential when conditional distributions require distinct optimal transport strategies that feature modulation approaches cannot capture. Importantly, we achieve this without introducing constraints on the neural network architectures (we demonstrate successful cases of MLPs and CNNs), achieving high generality and scope in terms of model design and application domain.
Experiments across scientific computing and image generation tasks show superior performance compared to alternative conditioning approaches. Systematic ablation studies confirm the value of each architectural component. For global sensitivity analysis, our framework replaces hundreds of individual OT problems with a single trained model, achieving high correlation ($\rho=0.942$) with traditional methods while enabling previously intractable analyses. For climate economic modeling, we efficiently emulate complex multivariate distributions under hybrid conditioning with preserved statistical properties.
Limitations and Future Work. Our evaluation focuses primarily on feedforward architectures with residual connections. Recurrent or attention-based architectures could potentially improve performance for sequential conditioning scenarios. While we handle discrete, continuous, and hybrid conditioning effectively, we have not explored complex modalities like CLIP embeddings or pixel-wise semantic labels that could enable more sophisticated generative modeling.
Our hypernetwork approach, though efficient relative to separate networks per condition, still costs more computationally than simple concatenation. More efficient hypernetwork designs or adaptive conditioning mechanisms could reduce this overhead. Theoretical analysis of convergence properties and approximation guarantees would strengthen the foundation for critical applications. Future directions include multi-modal conditioning, connections to diffusion models and normalizing flows, and applications to dynamical systems and causal inference. The modular framework design makes it well-suited for adaptation to emerging generative modeling paradigms and scientific computing applications.
@@
Broader Impacts. Our work addresses computational barriers in scientific computing and uncertainty quantification. Making conditional optimal transport computationally tractable enables more comprehensive sensitivity analysis for complex models in climate science, economics, and engineering—potentially improving policy decisions and scientific understanding. The approach also democratizes access to advanced optimal transport techniques for researchers with limited computational resources. For generative modeling, our framework enables more sophisticated conditional generation with principled transport-based approaches, complementing existing diffusion and flow-based methods. This could accelerate progress in controlled content generation, data augmentation, and domain adaptation. However, the ability to generate realistic samples conditioned on specific attributes could be misused for creating misleading synthetic media or other harmful content. Appropriate safeguards include access controls for sensitive applications, combining automated analysis with domain expertise for critical decisions, and providing educational resources to help practitioners understand limitations. The substantial benefits for scientific computing and uncertainty quantification outweigh potential risks when proper oversight is implemented.
Acknowledgments
Carlos Rodriguez-Pardo, Leonardo Chiani, and Massimo Tavoni acknowledge support from the European Research Council, ERC grant agreement number 101044703 (EUNICE) CUP D87G22000340006.